Perine, Inc., has balance sheet equity of $6.2 million. At the same time, the income statement shows net income of $948,600. The company paid dividends of $493,272 and has 100,000 shares of stock outstanding. If the benchmark PE ratio is 26, what is the target stock price in one year? Assume the firm will grow at the sustainable growth rate.
Answer #1

ROE = Net income / Equity

ROE = $948,600 / $6,200,000

ROE = .1530, or 15.30%

Retention ratio = 1 − Dividends / Net income

Retention ratio = 1 − $493,272 / $948,600

Retention ratio = .4800, or 48.00%

Sustainable growth rate = (ROE × b) / (1 − ROE × b)

Sustainable growth rate = (0.1530 × 0.4800) / (1 − 0.1530 × 0.4800)

Sustainable growth rate = .0793 or 7.93%

EPS0 = Net income / Shares outstanding

EPS0 = $948,600 / 100,000

EPS0 = $9.49

EPS1 = EPS0(1 + g)

EPS1 = $9.49(1 + .0793)

EPS1 = $10.24

P1 = Benchmark PE ratio × EPS1

P1 = 26($10.24)

P1 = $266.18